UITF Meeting - January 29, 2018
UITF Schedule - Zimbra calendar resource now functional. Add uitf-schedule to your zimbra calendar to see cae lock-up dates/times, etc.
- Choppers are being conditioned by Mark and Larry. Thanks!
- What is status of 1X/1Y and 2X/2Y?
- When do you estimate to complete conditioning for 200 keV beam (75 Watts RF power on choppers), at least 1X/2X
- We'll need screens/guidance to operate choppers - how's that going?
- Address PS rack ITF03 which prevents using four magnets presently
- Need to add one more corrector pair between the gun and the solenoid
- EPICS magnet screen is complete and functional with field maps provided by Joe
- Insulate and strain relieve the two FD chopper solenoids (they're off now, M. Beck is working on it)
- When will the two MDC valves have status read back?
- Performed software functionality test in UITF control room last week.
- What else is needed for full functionality?
- Cables installed and terminated.
- Status on hardware, firmware, and EPICS software?
- We need them operational and reading in EPICS before running beam in Tune Mode
- From last meeting: Cool down transfer line and 1/4 cryo-module to 80 K.
- Estimated time to complete?
- Update: D. Bullard completed installation of guard vacuum relief vent outside the cave
- UPDATES: cryogenic connections?
- Update on RF for Quarter Cryo Module?
- Estimate time to complete?
- UPDATES: Klystron connections?
Epics screens Working with T. Larrieu and team on SAVE/RESTORE.
Matt is working on OSP for buncher and choppers
By February 5, 2018:
All cryo work complete All High Power RF work complete OSP for QCM approved ODH assessment, vent can on relief stack, penetrations filled, etc,